A bakery finds that the price they can sell cakes is given by the function p = 580 − 10x where x is the number of cakes sold per day, and p is price. The total cost function of the company is given by c = (30+5x) 2 where x is previously defined, and c is total cost.
A. Find the revenue and marginal revenue functions [Hint: revenue is price multiplied by quantity i.e. revenue = price × quantity] .B. Find the fixed cost and marginal cost function [Hint: fixed cost does not change with quantity produced] . C. Find the profit function [Hint: profit is revenue minus total cost] . D. Find the quantity that maximizes profit.
1
Expert's answer
2020-05-04T14:59:39-0400

A.

Revenue function = (price per unit) "\times" (quantity of units)


"R=p(x)\cdot x=(580-10x)x=580x-10x^2"

Total point


"C(x)=(30+5x)^2=900+300x+25x^2"

Marginal revenue


"R'(x)={dR \over dx}=(580x-10x^2)'=580-20x"

B.


"C(x)=900+300x+25x^2"

"Fixed\ cost=900"

"Marginal\ cost=C'(x)=(900+300x+25x^2)'=""=300+50x"

C.

Profit function = revenue − cost


"P(X)=R(x)-C(x)=""=580x-10x^2-(900+300x+25x^2)=""=-35x^2+280x-900, x\geq0"

Marginal profit


"Marginal\ profit=P'(x)=""=(-35x^2+280x-900)'=-70x+280"

"P'(x)=0=>-70x+280=0=>x=4"

If "0\leq x<4," then "P'(x)>0, P(x)" increases.

If "x>4," then "P'(x)<0,P(x)" decreases.


"P(4)=-35(4)^2+280(4)-900=-340"

The function "P(x)" is the local maximum with value of "-340" at "x=4."

Since the function "P(x)" has the only extremum, then the profit function has the absolute maximum with value of "-340" at "x=4."